OpenStudy (peterg2001):

add -2/3n and 7/3. @recon14193 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

-2/3n+7/3 they connot be added as they are not like terms

OpenStudy (alexandervonhumboldt2):

you cannot add sheps with cabage it is similar

OpenStudy (peterg2001):

@StudyGurl14

OpenStudy (studygurl14):

\(\large -\frac{2}{3n}+\frac{7}{3}\rightarrow -\frac{2}{3n}+\frac{7n}{3n}\) now you can add. but i don't really see the point in doing this

OpenStudy (alexandervonhumboldt2):

well they mean in sipliest form i think

OpenStudy (alexandervonhumboldt2):

what are the choises?